A boy throws a ball vertically at a velocity of 4.8 m s−1.
(a) How long is it before he catches it again?
(b) What will be the ball’s greatest height above the point
of release?
1
Expert's answer
2020-11-20T06:53:43-0500

"V=u+at"


"-v=u+at"


"-4.8=4.8-gt"


"9t=9.8"


"t= 1 sec"


b) "v^{2}=u^{2}+2as"


"0=(4.8)^2-2(9.8)"


"S= \frac{4.8}{2\times 9.8}"


"S= 1.175m"